# High-Dimensional Sparse Clustering via Iterative Semidefinite Programming Relaxed K-Means

### v1

@@ Appendix G Supplementary details for numerical study
This section provides additional details on the numerical studies, omitted from Section 4. Appendix G.1 provides details on parameter tuning, early stopping criterion, implementation details, and computing resources. Appendix G.2 describes sparsity-aware high-dimensional clustering methods used for comparison. Appendix G.3 presents two additional simulation results: 1) numerical evidence supporting the convergence of our algorithm under the proposed stopping criterion, and 2) evaluation of Algorithm 2 under known non-isotropic covariance.
G.1 Supplementary details for the simulations
@@
This section provides details on hyperparameter tuning, software implementation, early stopping criteria, and computing resources related to the simulation study presented in Section 4.
Parameter tuning for lasso.
@@
For regularization parameter tuning in the ISEE subroutine (Algorithm 4) at each iteration of Algorithm 3, we use the lasso path algorithm to efficiently explore a range of regularization values and select the one that minimizes the Akaike Information Criterion (AIC). This procedure is implemented using functions from the MATLAB package penalized (McIlhagga,, 2016).
Early stopping.
@@
For efficiency, we employ a data-driven stopping criterion to terminate the algorithm before reaching the $T$th iteration. At each iteration $t$, the stopping condition is met if either of the following conditions is met: (1) the relative change in the objective value falls below $\pi\%$; or (2) after a warm-up period of $w$ iterations, the best objective value observed shows less than $\pi\%$ improvement over the most recent $\eta$ iterations. The algorithm terminates only when this condition holds for both the SDP objective in SDP K-means subroutine (line 1 of Algorithm 1) and the original K-means objective. Unless stated otherwise, we set $w=10$, $\eta=5$, and $\pi=1\%$ in our experiments. Appendix G.3 provides a numerical evaluation of this stopping criterion.
Software implementation.
@@
Our algorithms are implemented in MATLAB, and all simulations are run using version R2023b. To solve the SDP problems, we use the ADMM-based solver SDPNAL++ (Sun et al.,, 2020), which scales efficiently for sample sizes up to $n<5000$. In the final step of the SDP-relaxed K-means, we extract the leading eigenvector (an $n$-dimensional vector) from the output matrix $Z$ and apply the kmeans function from MATLAB Statistics and Machine Learning Toolbox. For the ISEE subroutine, we utilize the MATLAB Parallel Computing Toolbox to simulataeneously estimate the cluster centers and precision matrix blocks. For graphical lasso in Algorithm 5, we use the implementation provided by Leung et al., (2016). In Scenario 1, hierarchical clustering is performed using the linkage function with default options, also from MATLAB Statistics and Machine Learning Toolbox. When comparing against baseline methods, we use their original implementations, some in MATLAB and others in R, executed within their respective environments. All implementations and simulation codes, including those for the baseline methods written by their original authors, are available at our GitHub repository.
@@ G.2 Sparsity-aware clustering methods used for benchmarking
This section provides details of the baseline sparsity-aware methods that were compared to our algorithms in Section 4.
IF-PCA.
@@
Influential Feature PCA (IF-PCA; Jin and Wang, 2016) is a two step method. In the variable selection step, features with the highest Kolmogorov-Smirnov (KS) scores are selected, with the selection threshold determined in a data-driven manner using the Higher Criticism principle (Donoho and Jin,, 2004). In the clustering step, spectral clustering is applied: the top $K-1$ left singular vectors are extracted from the normalized data matrix after feature selection and used as inputs to the classical $K$-means algorithm to estimate cluster labels. For simulations, we use the authors’ MATLAB implementation.
SAS.
@@
Sparse Alternate Similarity clustering (SAS; Arias-Castro and Pu, 2017) is an iterative method that uses the fact that the within-cluster dissimilarity of K-means, measured by the squared $\ell_{2}$ distances, can be decomposed coordinate-wise. The algorithm alternates between clustering and support estimation: given a support set, it applies K-means to estimate the clusters; given the cluster assignments, it selects the top $\hat{|S_{0}|}$ coordinates with the largest average within-cluster dissimilarity. The tuning parameter $\hat{|S_{0}|}$, which denotes the number of signal features, is selected using the following permutation-based approach: For each $\hat{|S_{0}|}=1,\ldots,p$, run the SAS algorithm to obtain corresponding cluster and support estimates. For each run, compare the within-cluster dissimilarity of the selected features under the estimated cluster assignment to that under a randomly permuted cluster assignment. The discrepancy between the two is quantified using the gap statistic (Tibshirani et al.,, 2001). Then choose the value of $\hat{|S_{0}|}$ that yields the largest gap statistic. We use the authors’ R implementation.
Sparse K-means.
@@
Witten and Tibshirani, (2010) proposes Sparse K-means, an iterative method which exploits the fact that the within-cluster dissimilarity in K-means, measured by squared $\ell_{2}$ distances, can be decomposed coordinate-wise. The method assigns a feature weight $w_{j}$ to each variable $j=1,\ldots,p$, and constructs a feature-weighted version of the K-means objective. It then optimizes this objective over both feature weights and cluster assignments, subject to a combined lasso and ridge penalty. The proposed algorithm alternates between maximizing the objective with respect to the cluster assignments (given fixed feature weights) and with respect to the feature weights (given fixed cluster assignments). For simulations, we use the author’s R package sparcl.
CHIME.
@@
CHIME (Cai et al.,, 2019) is an EM-type iterative method, where the E-step updates the cluster assignments based on current Gaussian mixture model parameters, and the M-step updates the model parameters with cluster labels fixed. Notably, CHIME avoids explicit covariance estimation by employing an $\ell_{1}$-regularized estimation of the sparse discriminative direction $\boldsymbol{\beta}^{\ast}=\boldsymbol{\Omega}^{\ast}(\boldsymbol{\mu}_{1}^{\ast}-\boldsymbol{\mu}_{2}^{\ast})$. After a suitable number of iterations, the estimated $\boldsymbol{\beta}^{\ast}$ is plugged into the Fisher discriminant rule to produce the final cluster assignments. For simulations, we use the authors’ MATLAB implementation.
@@
6:     $\dot{\mathbf{X}}^{t}\leftarrow\bigl{(}\mathbf{X}_{\cdot,\hat{G}^{t-1}_{1}}-\hat{\bar{\mathbf{X}}}_{1}^{t}\mathbf{1}_{\hat{G}^{t-1}_{1}}^{\top},\;\mathbf{X}_{\cdot,\hat{G}^{t-1}_{2}}-\hat{\bar{\mathbf{X}}}_{2}^{t}\mathbf{1}_{\hat{G}^{t-1}_{2}}^{\top}\bigr{)}$triangleright Cluster-wise demeaned data matrix
@@
10:     $\hat{S}^{t}\leftarrow\left\{j\in\{1,\ldots,p\}:|\hat{\beta}^{t}_{j}|>\sqrt{2\hat{\omega}_{jj}^{t}n\log(2p)}/\sqrt{|\hat{G}^{t-1}_{1}|\cdot|\hat{G}^{t-1}_{2}|}\right\}$triangleright Variable selection
@@
13:     $\hat{G}_{1}^{t},\hat{G}_{2}^{t}\leftarrow\texttt{SDPkmeans}(\hat{\tilde{\mathbf{X}}}^{t}_{\hat{S}^{t},\cdot},\hat{\boldsymbol{\Sigma}}^{t}_{\hat{S}^{t},\hat{S}^{t}})$triangleright Clustering (Algorithm 1)
@@
Next, to assess the efficacy of Algorithm 3, which avoids full precision matrix estimation and uses a cutoff based on the convergence rate of ISEE, we introduce two modified variants.
Algorithm 5: Algorithm 3 with graphical lasso.
At each iteration $t$, Algorithm 5 fully estimates the precision matrix using graphical lasso (Friedman et al.,, 2008), and uses it to compute the necessary quantities: $\tilde{\mathbf{X}}$, $\boldsymbol{\beta}^{\ast}$, and $\boldsymbol{\Sigma}_{\hat{S}^{t},\hat{S}^{t}}$. These estimates are then used in variable selection and clustering steps.
Algorithm 6: Algorithm 3 with threshold based on Gaussian maximal inequality.
@@
6:     $\hat{S}^{t}\leftarrow\left\{j\in\{1,\ldots,p\}:|\hat{\beta}^{t}_{j}|>\sqrt{2\hat{\omega}_{jj}^{t}n\log(2p)}/\sqrt{|\hat{G}^{t-1}_{1}|\cdot|\hat{G}^{t-1}_{2}|}\right\}$triangleright Variable selection
@@
8:     $\hat{G}_{1}^{t},\hat{G}_{2}^{t}\leftarrow\texttt{SDPkmeans}(\hat{\tilde{\mathbf{X}}}^{t}_{\hat{S}^{t},\cdot},\;\hat{\boldsymbol{\Sigma}}^{t}_{\hat{S}^{t},\hat{S}^{t}})$triangleright Clustering (Algorithm 1)
@@
G.3 Additional simulation results
This section presents two additional simulation results.
G.3.1 Numerical Evidence Supporting the Convergence of Our Algorithm
@@
We provide numerical evidence supporting the convergence of Algorithm 3 under the stopping criterion presented in Appendix G.1. We re-use the setup from Scenario 2 of Section 4, speicifically the scenario $(p,\Delta_{\boldsymbol{\Sigma}^{\ast}},\rho)=(400,4,0.45)$. Line 1 of Table 3 presents the details of the setting. The stopping criteria are applied using the hyperparameters $T=100$, $\pi=0.5\%$, $w=40$, and $\eta=20$. We conduct 200 independent replications to record when the iterative algorithm 3 terminates. At each stopping point, we record the mis-clustering rate, the number of true positives (signal entries retained), and false positives (noise entries retained). The results in Figure 3 reveal two main insights about the behavior of our stopping criteria and the performance of Algorithm 3: First, the stopping criteria are effective. Figure 3-(a) shows that the algorithms consistently stop before reaching the maximum iteration limit, providing empirical evidence of convergence. Across different estimation-cutoff strategies, the stopping points are similarly concentrated. Second, true discoveries contribute more to clustering performance than the avoidance of false ones. As shown in Figure 3-(b), the original Algorithm 3 demonstrates more stable clustering performance compared to its variant, which often yields mis-clustering rates near 0.5. Figures 3-(c) and 3-(d) shows that the original algorithm makes slightly more true discoveries and incurs many more false discoveries, respectively.

### v2

@@ Appendix G DETAILS FOR NUMERICAL STUDY
This section provides additional details on the numerical studies, omitted from Section 4. Appendix G.1 provides details on parameter tuning, early stopping criterion, implementation details, and computing resources. Appendix G.2 describes sparsity-aware high-dimensional clustering methods used for comparison.
G.1 Supplementary details for the simulation studies
@@
This section provides details on hyperparameter tuning, software implementation, early stopping criteria, and computing resources related to the simulation studies presented in Section 4 and Appendix H.
Initialization
Unless specified, we use spectral clustering (Han et al.,, 2023) for initialization of our iterative algorithms.
High-dimensional regression
@@
We use the lasso (Tibshirani,, 1996) for high-dimensional regression between nodes in the ISEE subroutine (Algorithm 4) of Algorithm 3. At each iteration, the regularization parameter is tuned by computing the lasso path and selecting the value that minimizes the Akaike Information Criterion (AIC). This procedure is implemented using functions from the MATLAB package penalized (McIlhagga,, 2016).
Early stopping
@@
For efficiency, we employ a data-driven stopping criterion to terminate the algorithm before reaching the $T$th iteration. At each iteration $t$, the stopping condition is met if either of the following conditions is met: (1) the relative change in the objective value falls below $\pi\%$; or (2) after a warm-up period of $w$ iterations, the best objective value observed shows less than $\pi\%$ improvement over the most recent $\eta$ iterations. The algorithm terminates only when this condition holds for both the SDP objective in SDP $K$-means subroutine (line 1 of Algorithm 1) and the original $K$-means objective. Unless stated otherwise, we set $w=10$, $\eta=5$, and $\pi=1\%$ in our experiments. Appendix H provides a numerical evaluation of this stopping criterion.
Software implementation
@@
Our algorithms are implemented in MATLAB, and all simulations are run using version R2023b. To solve the SDP problems, we use the ADMM-based solver SDPNAL++ (Sun et al.,, 2020). In the final step of the SDP-relaxed $K$-means, we extract the leading eigenvector (an $n$-dimensional vector) from the output matrix $Z$ and apply the kmeans function from MATLAB Statistics and Machine Learning Toolbox. For the ISEE subroutine, we utilize the MATLAB Parallel Computing Toolbox to simulataeneously estimate the cluster centers and precision matrix blocks. For graphical lasso in Algorithm 5, we use the implementation provided by Leung et al., (2016). In Scenario 1, hierarchical clustering is performed using the linkage function with default options, also from MATLAB Statistics and Machine Learning Toolbox. When comparing against baseline methods, we use their original implementations, some in MATLAB and others in R, executed within their respective environments. All implementations and simulation codes, including those for the baseline methods written by their original authors, are available at our GitHub repository.
@@ G.2 Sparsity-aware clustering methods used for benchmarking
This section describes the baseline sparsity-aware methods used for comparison with our algorithms in the numerical studies presented in Section 4.
@@
IFPCA  Influential Feature PCA (IFPCA; Jin and Wang,, 2016) is a two step method. In the feature selection step, features with the highest Kolmogorov-Smirnov (KS) scores are selected, with the selection threshold determined in a data-driven manner using the Higher Criticism principle (Donoho and Jin,, 2004). In the clustering step, spectral clustering is applied: the top K1K-1 left singular vectors are extracted from the normalized data matrix after feature selection and used as inputs to the classical $K$-means algorithm to estimate cluster labels. For simulations, we use the authors’ MATLAB implementation.
@@
SAS  Sparse Alternate Similarity clustering (SAS; Arias-Castro and Pu,, 2017) is an iterative method that uses the fact that the within-cluster dissimilarity of $K$-means, measured by the squared $\ell_{2}$ distances, can be decomposed coordinate-wise. The algorithm alternates between clustering and support estimation: given a support set, it applies $K$-means to estimate the clusters; given the cluster assignments, it selects the top $\hat{|S_{0}|}$ coordinates with the largest average within-cluster dissimilarity. The tuning parameter $\hat{|S_{0}|}$, which denotes the number of signal features, is selected using the following permutation-based approach: For each $\hat{|S_{0}|}=1,\ldots,p$, run the SAS algorithm to obtain corresponding cluster and support estimates. For each run, compare the within-cluster dissimilarity of the selected features under the estimated cluster assignment to that under a randomly permuted cluster assignment. The discrepancy between the two is quantified using the gap statistic (Tibshirani et al.,, 2001). Then choose the value of $\hat{|S_{0}|}$ that yields the largest gap statistic. We use the authors’ R implementation.
@@
SKM  Sparse $K$-means (SKM; Witten and Tibshirani,, 2010 ) is an iterative method which exploits the fact that the within-cluster dissimilarity in $K$-means, measured by squared $\ell_{2}$ distances, can be decomposed coordinate-wise. The method assigns a feature weight $w_{j}$ to each variable $j=1,\ldots,p$, and constructs a feature-weighted version of the $K$-means objective. It then optimizes this objective over both feature weights and cluster assignments, subject to a combined lasso and ridge penalty. The proposed algorithm alternates between maximizing the objective with respect to the cluster assignments (given fixed feature weights) and with respect to the feature weights (given fixed cluster assignments). For simulations, we use the authors’ R package sparcl.
@@
CHIME CHIME (Cai et al.,, 2019) is an EM-type iterative method, where the E-step updates the cluster assignments based on current Gaussian mixture model parameters, and the M-step updates the model parameters with cluster labels fixed. Notably, CHIME avoids explicit covariance estimation by employing an $\ell_{1}$-regularized estimation of the sparse discriminative direction $\boldsymbol{\beta}^{\ast}=\boldsymbol{\Omega}^{\ast}(\boldsymbol{\mu}_{1}^{\ast}-\boldsymbol{\mu}_{2}^{\ast})$. After a suitable number of iterations, the estimated $\boldsymbol{\beta}^{\ast}$ is plugged into the Fisher discriminant rule to produce the final cluster assignments. For simulations, we use the authors’ MATLAB implementation. This algorithm requires careful initialization of the model parameters. For simulations, we used the true values of model parameters with very small Gaussian noise (standard deviation 0.01) added to each entry as the initial values.
Appendix H ADDITIONAL NUMERICAL RESULTS
This section supplements Section 4 by presenting additional simulation results.
H.1 Evaluation of Algorithm 2 under known non-isotropic covariance
As argued in Section 3.1 and further justified in Appendix E.2, the SDP problem in line 1 of Algorithm 1:
$\arg\max_{\mathbf{Z}\in\mathbb{R}^{n\times n}}$~ $\langle\tilde{\mathbf{X}}_{\hat{S}^{t},\cdot}^{\top}\boldsymbol{\Sigma}_{\hat{S}^{t},\hat{S}^{t}}\tilde{\mathbf{X}}_{\hat{S}^{t},\cdot},\mathbf{Z}\rangle$
s.t. $\mathbf{Z}^{\top}=~\mathbf{Z},~\mathbf{Z}\succeq~0,~\mathrm{tr}(\mathbf{Z})=~2,\mathbf{Z}\boldsymbol{1}_{n}=~\boldsymbol{1}_{n}$
captures the key sparse clustering signal encoded in $\boldsymbol{\beta}^{\ast}$. To validate this claim, we evaluate Algorithm 2 under the same setting as Figure 1 (line 1 of Table 1), but with a non-isotropic covariance. In particular, we adopt an AR(1) covariance structure. The data are generated under the model (16), with parameter settings folloiwng the line 1 of Table 3. The result presented in Figure 3 highlights two key findings. First, Algorithm 2 successfully adapts to the sparsity of $\boldsymbol{\beta}^{\ast}$: its performance remains stable, particularly when $\Delta_{\boldsymbol{\Sigma}^{\ast}}=4$, as sparsity increases while the total clustering signal is held fixed. Second, when the covariance structure is known, stronger correlations improve clustering. As $\rho$ increases, mis-clustering rates deteriorate more slowly with increasing sparsity while the total clustering signal is held fixed. This is expected, since knowing a high-correlation structure is effectively equivalent to working in a lower-dimensional space.
Figure 3: Mis-clustering rates of Algorithm 2 with spectral initialization under AR(1) correlation structures with $\rho=0.2,0.5,0.8$. The data generation follows line 1 of Table 3. Dotted lines of the same color indicate the mis-clustering rates of spectral initialization under the corresponding covariance structure.
Table 3: Summary of Parameter Settings for Model (16) used in Appendices H.1 - H.3, with $i,j=1,\ldots,p$.
Section Methods Evaluated Covariance structure $p$ $n$ Separation
H.1
Algorithm 2
(for known $\boldsymbol{\Sigma}^{\ast}$)
known, $\Sigma^{\ast}_{i,j}=\rho^{|i-j|}$,
$\rho\in\{0.2,0.5,0.8\}$
50-5000 200 $\Delta_{\boldsymbol{\Sigma}^{\ast}}^{2}=3^{2},4^{2}$
H.2
Algorithm 3
and its variants
unknown, $\Omega^{*}_{i,i+1}=\Omega^{*}_{i+1,i}=\rho$,
$\Omega^{*}_{i,i}=1$, $\Omega^{*}_{i,j}=0$ o.w.,
$\rho\in\{0.2,0.45\}$
400 500 $\Delta_{\boldsymbol{\Sigma}^{\ast}}^{2}=3^{2},4^{2}$
H.3
Early stopping of
Algorithm 3
unknown, $\Omega^{*}_{i,i+1}=\Omega^{*}_{i+1,i}=\rho$,
$\Omega^{*}_{i,i}=1$, $\Omega^{*}_{i,j}=0$ o.w., $\rho=0.45$
400 500 $\Delta_{\boldsymbol{\Sigma}^{\ast}}^{2}=4^{2}$
H.2 Impact of Parameter Estimation and Threshold Choice on Algorithm 3
We argued in Section 3.2 that the strength of Algorithm 3 lies in avoiding full estimation of the precision matrix. On the other hand, its threshold choice is based on the theoretical gaurantee of ISEE (Fan and Lv,, 2016), which involves many constants (we suppressed this by $\sqrt{\log n}$). Recall that for the threshold in Algorithm 2, only a single constant $\delta$ arises from the family-wise error rate derived by maximal inequality for Gaussians, which was also conveniently set to a fixed constant without theoretical background. Nevertheless, we argue that these constant factors have little practical impact, since the threshold can be set slightly conservative without loss of performance. This is justified by Theorem 1, which shows that SDP $K$-means is robust to overestimation of $S_{0}$. In this section, we numerically validate these points by comparing Algorithm 3 with two specially designed variants.
Algorithm 5: Algorithm 3 with graphical lasso
At each iteration $t$, Algorithm 5 fully estimates the precision matrix using graphical lasso (GraphicalLasso; Friedman et al.,, 2008), and uses it to compute the necessary quantities: $\tilde{\mathbf{X}}$, $\boldsymbol{\beta}^{\ast}$, and $\boldsymbol{\Sigma}_{\hat{S}^{t},\hat{S}^{t}}$. These estimates are then used in feature selection and clustering steps.
@@ H.2 Impact of Parameter Estimation and Threshold Choice on Algorithm 3
6:  $\dot{\mathbf{X}}^{t}\leftarrow\bigl(\mathbf{X}_{\cdot,\hat{G}^{t-1}_{1}}-\hat{\bar{\mathbf{X}}}_{1}^{t}\mathbf{1}_{\hat{G}^{t-1}_{1}}^{\top},\;\mathbf{X}_{\cdot,\hat{G}^{t-1}_{2}}-\hat{\bar{\mathbf{X}}}_{2}^{t}\mathbf{1}_{\hat{G}^{t-1}_{2}}^{\top}\bigr)$ $\triangleright$ Cluster-wise demeaned data matrix
@@
10:  $\hat{S}^{t}\leftarrow\left\{j\in\{1,\ldots,p\}:|\hat{\beta}^{t}_{j}|>\sqrt{2\hat{\omega}_{jj}^{t}n\log(2p)}/\sqrt{|\hat{G}^{t-1}_{1}|\cdot|\hat{G}^{t-1}_{2}|}\right\}$ $\triangleright$ feature selection
@@
13:  $\hat{G}_{1}^{t},\hat{G}_{2}^{t}\leftarrow\texttt{clust}(\hat{\tilde{\mathbf{X}}}^{t}_{\hat{S}^{t},\cdot},\hat{\boldsymbol{\Sigma}}^{t}_{\hat{S}^{t},\hat{S}^{t}})$ $\triangleright$ Clustering (Algorithm 1)
@@
Algorithm 6: Algorithm 3 with threshold based on Gaussian maximal inequality
@@
6:  $\hat{S}^{t}\leftarrow\left\{j\in\{1,\ldots,p\}:|\hat{\beta}^{t}_{j}|>\sqrt{2\hat{\omega}_{jj}^{t}n\log(2p)}/\sqrt{|\hat{G}^{t-1}_{1}|\cdot|\hat{G}^{t-1}_{2}|}\right\}$ $\triangleright$ feature selection
@@
8:  $\hat{G}_{1}^{t},\hat{G}_{2}^{t}\leftarrow\texttt{clust}(\hat{\tilde{\mathbf{X}}}^{t}_{\hat{S}^{t},\cdot},\;\hat{\boldsymbol{\Sigma}}^{t}_{\hat{S}^{t},\hat{S}^{t}})$ $\triangleright$ Clustering (Algorithm 1)
@@
Figure 4 compares Algorithm 3 with the two variants under the setting specified in line 2 of Table 3. Figure 4 highlights two key findings. First, Algorithm 5 performs poorly compared to Algorithm 3, indicating that avoiding full estimation of the precision matrix improves performance. Second, Algorithm 6 performs slightly worse than Algorithm 3 in some scenarios, suggesting that an exact feature selection cutoff is not necessary and reconfirming the robustness of SDP $K$-means to slight overestimation of $S_{0}$.
Figure 4: Mis-clustering rates of Algorithm 3, its two variants, and SAS (the best-performing baseline method in Figure 2). The settings follow the unknown covariance scenario with varying conditional correlations, as specified in line 2 of Table 3.
H.3 Numerical Evidence Supporting the Convergence of Our Algorithm
@@
We provide numerical evidence supporting the convergence of Algorithm 3 under the stopping criterion in Appendix G.1, and reaffirm that slight overestimation of $S_{0}$ is not harmful. We reuse one setting from Appendix H.2, specifically the scenario $(p,\Delta_{\boldsymbol{\Sigma}^{\ast}},\rho)=(400,4,0.45)$. Line 3 of Table 3 provides the details of this setting. The stopping criteria are applied using the hyperparameters $T=100$, $\pi=0.5\%$, $w=40$, and $\eta=20$. We conduct 200 independent replications to record when Algorithm 3 terminates its iteration. At each stopping point, we record the mis-clustering rate, the number of true positives (signal entries retained), and false positives (noise entries retained). The results in Figure 5 reveal two main insights about the behavior of our stopping criteria and the performance of Algorithm 3: First, the stopping criteria are effective. Figure 5-(a) shows that the algorithms consistently stop before reaching the maximum iteration limit, providing empirical evidence of convergence. Across different estimation-cutoff strategies, the stopping points are similarly concentrated. Second, true discoveries contribute more to clustering performance than the avoidance of false ones. As shown in Figure 5-(b), the original Algorithm 3 demonstrates more stable clustering performance compared to its variant, which often yields mis-clustering rates near 0.5. Figures 5-(c) and 5-(d) shows that Algorithm 3 incurs many more false discoveries. Since Figure 4 shows that Algorithm 3 slightly outperforms Algorithm 6, these plots reaffirm that modest over-selection of features is not particularly harmful.